### v1

# Higher-order Sobolev and Rellich inequalities via iterated Talenti’s principle

Abstract.
In this paper we establish higher-order Sobolev and Rellich-type inequalities on non-compact Riemannian manifolds supporting an isoperimetric inequality. We highlight two notable settings: manifolds with non-negative Ricci curvature and having Euclidean volume growth (supporting Brendle’s isoperimetric inequality) and manifolds with non-positive sectional curvature (satisfying the Cartan–Hadamard conjecture or supporting Croke’s isoperimetric inequality). Our proofs rely on various symmetrization techniques, the key ingredient is an iterated Talenti’s comparison principle. The non-iterated version is analogous to the main result of Chen and Li [J. Geom. Anal., 2023].
Key words and phrases:
Iterated Talenti’s principle, Sobolev and Rellich inequalities, Riemannian manifolds
2020 Mathematics Subject Classification:
58J60, 26D10, 53C21, 46E35
C. Farkas and S. Kajántó are supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
1. Introduction
In the last century, Schwarz symmetrization became a fundamental tool in proving functional inequalities. The symmetric rearrangement of a set $\Omega\subset\mathbb{R}^{n}$, typically denoted by $\Omega^{\star}$, is a ball centered at the origin with the property $\operatorname{vol}(\Omega)=\operatorname{vol}(\Omega^{\star})$. In a similar manner, the symmetric rearrangement of a function $f\colon\Omega\to\mathbb{R}$ is a function $f^{\star}\colon\Omega^{\star}\to[0,\infty)$, whose level sets are the symmetrizations of the level sets of $|f|$.
When proving functional inequalities, one typically exploits the following two fundamental properties of symmetrizations (see e.g., Lieb and Loss [17]): the Cavalieri principle (for every $q>0$, the $L^{q}$-norm is preserved under symmetrization) and the Pólya–Szegő inequality (for every $p>1$, the $p$-Dirichlet energy decreases under symmetrization). If in addition, the given functional inequality involves (singular) weights, the toolkit extends by the Hardy–Littlewood inequality (the $L^{1}$-norm of the product of two functions increases under symmetrization). For more details we refer to §2.
An alternative way to investigate a functional inequality, is to analyze the associated partial differential equation. According to Talenti’s principle [20], if $u$ and $v$ are weak solutions of the problems
$\begin{cases}-\Delta u=f,&\text{in }\Omega,\\ u=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}-\Delta v=f^{\star},&\text{in }\Omega^{\star},\\ v=0,&\text{on }\partial\Omega^{\star},\end{cases}$
respectively, then almost everywhere in $\Omega^{\star}$ one has $u^{\star}\leq v$. By an iterative application of this result, in the spirit of Gazzola, Grunau, and Sweers [11], one can prove higher order functional inequalities, as well.
The common key ingredient in establishing the aforementioned results, is the following sharp isoperimetric inequality: If $\Omega\subset\mathbb{R}^{n}$ is a bounded set, then
$\operatorname{per}(\partial\Omega)\geq n\omega_{n}^{\frac{1}{n}}\cdot\operatorname{vol}(\Omega)^{\frac{n-1}{n}},$
where $\omega_{n}$ is the volume of the $n$-dimensional unit ball, while $\operatorname{per}(\cdot)$ and $\operatorname{vol}(\cdot)$ stands for the perimeter and volume, respectively. Moreover, equality holds if and only if $\Omega$ is a ball. In other words, the ball has the least perimeter among the sets with given volume.
The goal of the paper is to show that if an isoperimetric inequality holds on a Riemannian manifold, then various Euclidean functional inequalities extend naturally to that setting. Throughout the paper, let $(M,g)$ be an $n$-dimensional Riemannian manifold, and suppose that there exists a constant $c_{g}>0$ such that for every bounded $\Omega\subset M$ one has
$\operatorname{per}(\partial\Omega)\geq c_{g}\cdot n\omega_{n}^{\frac{1}{n}}\cdot\operatorname{vol}(\Omega)^{\frac{n-1}{n}}.$ (1)
Note that this assumption is reasonable according to Brendle’s isoperimetric inequality [5] on non-negatively curved spaces and to Croke’s isoperimetric inequality [8] on non-positively curved spaces, see §2.1.
In the sequel, assuming inequality (1), we present our main results. Hereafter, for every $m\geq 1$ let
$|D_{g}^{m}u|\stackrel{{\scriptstyle\textup{def}}}{{=}}\begin{cases}|\Delta_{g}^{k}u|,&\text{if }m=2k,\\ |\nabla_{g}(\Delta_{g}^{k}u)|,&\text{if }m=2k+1.\end{cases}$
Our first result is a higher-order Sobolev inequality, which can be stated as follows.
Theorem 1.1.
Let $(M,g)$ be a complete, non-compact $n$-dimensional Riemannian manifold supporting (1). Let $m$ be a positive integer and $p>1$ such that $n>mp$. Then for every $u\in C_{0}^{\infty}(M)$ one has
$\int_{M}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}\geq(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{M}|u|^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$ (2)
where $p^{*}\stackrel{{\scriptstyle\textup{def}}}{{=}}\frac{np}{n-mp}$, and $S_{m,p}$ is the Euclidean Sobolev constant of degree $m$ and order $p$.
Note that $S_{m,p}$ is the greatest constant for which inequality (2) holds when $M=\mathbb{R}^{n}$ (and $c_{g}=1$). If $p=2$, according to Cotsiolis and Tavoularis [7], its exact value is given by
$S_{m,2}=\left(\pi^{m}n(n-2m)\prod_{s=1}^{m-1}(n^{2}-4s^{2})\right)^{-1}\left(\frac{\Gamma(n)}{\Gamma(\frac{n}{2})}\right)^{\frac{2m}{n}}.$
Our second result is the following higher-order Rellich inequality:
Theorem 1.2.
Let $(M,g)$ be a complete, non-compact $n$-dimensional Riemannian manifold supporting (1). Let $m$ be a positive integer and $p>1$ such that $n>mp$. Fix $x_{0}\in M$ and let $\rho(\cdot)=d_{g}(x_{0},\cdot)$ denote the Riemannian distance from $x_{0}$. Then for every $u\in C_{0}^{\infty}(M)$ one has
$\int_{M}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}\geq(c_{g})^{mp}\cdot R_{m,p}\cdot\int_{M}\frac{|u|^{p}}{\rho^{mp}}\,\mathrm{d}v_{g},$ (3)
where $R_{m,p}$ is the Euclidean Rellich constant of degree $m$ and order $p$.
Note that $R_{m,p}$ is the greatest constant for which inequality (3) hold when $M=\mathbb{R}^{n}$ (and $c_{g}=1$). According to Mitidieri [19], its exact value is given by
$R_{m,p}=\begin{cases}\displaystyle\prod_{s=1}^{k}\left(\frac{n}{p}-2s\right)^{p}\left(\frac{n(p-1)}{p}+2s-2\right)^{p},&\text{if }m=2k,\\ \displaystyle\left(\frac{n-p}{p}\right)^{p}\prod_{s=1}^{k}\left(\frac{n}{p}-2s-1\right)^{p}\left(\frac{n(p-1)}{p}+2s-1\right)^{p},&\text{if }m=2k+1.\end{cases}$
Our third result is an extension of the previous theorem, which allows for discussing ‘improved’ Rellich inequalities.
Theorem 1.3.
Let $(M,g)$ be a complete, non-compact $n$-dimensional Riemannian manifold supporting (1). Let $m$ be a positive integer and $p>1$ such that $n>mp$. Fix $\Omega\subset M$, $x_{0}\in\Omega$ and denote by $\rho(\cdot)=d_{g}(x_{0},\cdot)$ the Riemannian distance from $x_{0}$. Suppose that $h\colon[0,\infty)\to[0,\infty)$ is a non-increasing function. Then for every $u\in C_{0}^{\infty}(\Omega)$ the following inequality holds:
$\int_{\Omega}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}\geq(c_{g})^{mp}\cdot C_{m,p}\cdot\int_{\Omega}|u|^{p}h(\rho)\,\mathrm{d}v_{g},$
provided that
$\int_{\Omega^{\star}}|D_{g_{0}}^{m}u^{\star}|^{p}\,\mathrm{d}v_{g_{0}}\geq C_{m,p}\cdot\int_{\Omega^{\star}}|u^{\star}|^{p}h(|x|)\,\mathrm{d}v_{g_{0}}.$
The main technical ingredient underlying the above results is an iterated Talenti’s principle, established on Riemannian manifolds verifying the isoperimetric inequality (1), see Theorem 2.3. The proof of the non-iterated version (see Theorem 2.2) is analogous to the main result of Chen and Li [6], which provides a Talenti’s principle on Riemannian manifolds having $\mathbf{Ric}_{g}\geq 0$ and Euclidean volume growth.
The paper is structured as follows. In §2 we present an iterated Talenti’s principle. First, in §2.1 we recall various isoperimetric inequalities. Next, in §2.2 we display the definition and relevant properties of Schwarz symmetrization in the Riemannian setting. Finally, in §2.3 we formulate the principle together with its iterated version (see Theorem 2.3). In §3 we prove Theorem 1.1, that is we establish higher-order Sobolev inequalities on Riemannian manifolds supporting the inequality (1). Moreover, in Proposition 3.1 in the particular case of non-positive Ricci curvature, Euclidean volume growth, and $m=p=2$, we present a sharpness result. In §4 we prove Theorem 1.2, that is we establish higher-order Rellich inequalities on Riemannian manifolds supporting the inequality (1). Here we also prove Theorem 1.3, which allows for establishing improved higher-order Rellich-type inequalities. We conclude the paper with various applications of Theorem 1.3.
2. An Iterated Talenti’s principle
In this section we present an iterated version of Talenti’s principle on Riemannian manifolds supporting the isoperimetric inequality (1). First, in §2.1 we recall some well-known Riemannian isoperimetric inequalities. Next, in §2.2 we present various concepts concerning the Schwarz symmetrization. Finally, in §2.3 we present an iterated Riemannian version of Talenti’s principle.
2.1. Isoperimetric inequalities
Throughout the paper we consider Riemannian manifolds $(M,g)$ that supports the isoperimetric inequality (1). To motivate this assumption, in what follows, we recall some well-known isoperimetric inequalities (together with their relevant properties) in the Riemannian setting.
Case 1: $\mathbf{Ric}_{g}\geq 0$. The asymptotic volume ratio of $(M,g)$ is given by
$\mathsf{AVR}_{g}\stackrel{{\scriptstyle\textup{def}}}{{=}}\lim_{r\to\infty}\frac{\operatorname{vol}_{g}(B_{g}(x,r))}{\omega_{n}r^{n}},$
where $\operatorname{vol}_{g}$ stands for the Riemannian volume induced by $g$; $B_{g}(x,r)$ denotes a geodesic ball in $M$ centered at $x\in M$ having radius $r$; while $\omega_{n}$ denotes the volume of the $n$-dimensional Euclidean unit ball. On the one hand, we note that the above limit is independent of the choice of $x\in M$. On the other hand, according to the Bishop–Gromov comparison theorem (see e.g., Gallot, Hulin, and Lafontaine [9, Theorem 3.101]), one has $\mathsf{AVR}_{g}\in[0,1]$. The typical assumption in this setting is that the manifold $(M,g)$ has Euclidean volume growth, that is
$\mathsf{AVR}_{g}>0.$
Under this assumption the following sharp isoperimetric inequality holds (see Brendle [5, Corollary 1.3]): For every bounded domain $\Omega\subset M$ one has
$\operatorname{per}_{g}(\partial\Omega)\geq\mathsf{AVR}_{g}^{\frac{1}{n}}\cdot n\omega_{n}^{\frac{1}{n}}\cdot\operatorname{vol}_{g}(\Omega)^{\frac{n-1}{n}},$ (4)
where $\operatorname{per}_{g}(\partial\Omega)$ and $\operatorname{vol}_{g}(\Omega)$ denote the perimeter and volume of $\Omega$, respectively. Moreover, equality occurs if and only if $(M,g)$ is isometric to the Euclidean space and $\Omega$ is isometric to an Euclidean ball. For an alternative proof (using optimal mass transport theory) of the above result see Balogh and Kristály [2].
Note that in case of the standard Euclidean space, denoted in the sequel by $(\mathbb{R}^{n},g_{0})$, one has $\mathsf{AVR}_{g_{0}}=1$, and thus, inequality (4) reduces to the Euclidean isoperimetric inequality. We also notice that inequality (4) is a particular case of inequality (1) for $c_{g}=\mathsf{AVR}^{1/n}$.
Case 2: $\mathbf{K}_{g}\leq 0$. This case is strongly related to the so called Cartan–Hadamard conjecture, which can be stated as follows. If $(M,g)$ is a Cartan–Hadamard manifold, that is a complete, simply connected Riemannian manifold with non-positive sectional curvature, then it supports the Euclidean isoperimetric inequality (i.e., inequality (1) with constant $c_{g}=1$). The conjecture is proved for $n\in\{2,3,4\}$, see Bol [4], Kleiner [15], and Croke [8], for higher dimensions it is still open. Fortunately, according to Croke the isoperimetric inequality (1) holds for $n\geq 3$ with the constant
$c_{g}=n^{-\frac{1}{n}}\omega_{n}^{1-\frac{2}{n}}\left((n-1)\omega_{n-1}\int_{0}^{\frac{\pi}{2}}\cos^{\frac{n}{n-2}}(t)\sin^{n-2}(t)\,\mathrm{d}t\right)^{\frac{2}{n}-1}.$ (5)
Note that $c_{g}\leq 1$, for every $n\geq 3$, while $c_{g}=1$ if and only if $n=4$.
Case 3: $\mathbf{Ric}_{g}\geq k\in\mathbb{R}$ and positive lower bound on volume of unit balls. According to Hebey [13, Lemma 3.2], suppose that $(M,g)$ is a complete $n$-dimensional Riemannian manifold with $\mathbf{Ric}_{g}\geq k$ for some $k\in\mathbb{R}$, and there exists $v>0$ such that
$\operatorname{vol}_{g}(B_{g}(x,1))\geq v,\quad\forall x\in M.$
Then there exist $c_{g},\eta>0$ depending only on $n$, $k$, and $v$ such that for every open $\Omega\subset M$ with smooth boundary, compact closure, and $\operatorname{vol}_{g}(\Omega)\leq\eta$, the isoperimetric inequality (1) holds.
Case 4: Bounded elliptic Kato constant As in Impera, Rimoldi, and Veronelli [14, Theorem 1.1], let $(M,g)$ be a complete non-parabolic Riemannian manifold with
$\limsup_{r\to\infty}\frac{\operatorname{vol}_{g}(B_{g}(y,r))}{r^{n}}\geq\beta>0.$
Suppose that the elliptic Kato constant satisfies $k_{\infty}<\frac{1}{n-2}$ and $\Omega\subset M$ is a compact domain with smooth boundary. Then isoperimetric inequality (1) holds with
$c_{g}=(1-(n-2)k_{\infty})^{\frac{4(n-1)}{n(n-2)}}\beta^{\frac{1}{n}}\omega_{n}^{-\frac{1}{n}}.$
2.2. Schwarz symmetrization
It turns out that the isoperimetric inequality (1) enables us to compare various quantities on a general Riemannian manifold to the corresponding quantities on the Euclidean space. This is precisely done using the symmetrization procedure presented below.
If $\Omega\subset M$ is a bounded domain, its symmetrization, denoted by $\Omega^{\star}$, is an Euclidean ball, centered at the origin satisfying
$\operatorname{vol}_{g}(\Omega)=\operatorname{vol}_{g_{0}}(\Omega^{\star}).$ (6)
Using the Euclidean isoperimetric inequality (with $c_{g_{0}}=1$), inequality (1) can be rewritten as
$\operatorname{per}_{g}(\partial\Omega)\geq c_{g}\cdot\operatorname{per}_{g_{0}}(\partial\Omega^{\star}).$ (7)
The Schwarz symmetrization of a function $u\colon\Omega\to\mathbb{R}$, is a function $u^{\star}\colon\Omega^{\star}\to[0,\infty)$, which satisfies
$\operatorname{vol}_{g}(\{x\in M:|u(x)|>t\})=\operatorname{vol}_{g_{0}}(\{x\in\mathbb{R}^{n}:u^{\star}(x)>t\}),\quad\forall t>0.$ (8)
In the sequel, we present three fundamental properties of symmetrized functions, which are exploited in the rest of the paper. First, the Cavalieri principle (see e.g. Lieb and Loss [17, Theorem 3.3/(iv)]) states that if $u\in L^{q}(\Omega)$ for some $q\in(0,\infty)$, then
$\|u\|_{L^{q}(\Omega)}=\|u^{\star}\|_{L^{q}(\Omega^{\star})}.$ (9)
Next, according to the Pólya–Szegő inequality (see e.g. Balogh and Kristály [2, Proposition 3.1]), if $|\nabla_{g}u|\in L^{p}(\Omega)$ for some $p>1$ and some smooth function $u$, then one has
$\|\nabla_{g}u\|_{L^{p}(\Omega)}\geq c_{g}\cdot\|\nabla_{g_{0}}u^{\star}\|_{L^{p}(\Omega^{\star})},$ (10)
where $\nabla_{g}$ and $\nabla_{g_{0}}$ denote the Riemannian and Euclidean gradient, respectively.
Finally, due to the Hardy–Littlewood inequality (see e.g. Lieb and Loss [17, Theorem 3.4]), if $x_{0}\in\Omega$ and $\rho(x)=d_{g}(x_{0},x)$, $u\in L^{p}(\Omega)$ for some $p>1$, and $h\colon[0,\infty)\to[0,\infty)$ is a non-increasing function, then
$\int_{\Omega}u(x)^{p}h(\rho(x))\,\mathrm{d}v_{g}\leq\int_{\Omega^{\star}}u^{\star}(x)^{p}h(|x|)\,\mathrm{d}v_{g_{0}}.$ (11)
Remark 2.1.
We note that some authors condiser an alternative symmetrization (tipically denoted by ). In this case, $\Omega^{\sharp}$ is an Euclidean ball satisfying $\operatorname{vol}_{g}(\Omega)=(c_{g})^{n}\cdot\operatorname{vol}_{g_{0}}(\Omega^{\sharp})$ instead of (6), and inequality (7) becomes $\operatorname{per}_{g}(\partial\Omega)\geq(c_{g})^{n}\cdot\operatorname{per}_{g_{0}}(\partial\Omega^{\sharp})$. At first sight, the same exponent of $c_{g}$ in the latter relations may seem to have computational advantage. However, when proving the alternative versions of the fundamental properties (9), (10) & (11), one needs to compare sets with the same volume, and then to perform changes of variables of the form $y=c_{g}\cdot x$. Using the $\Omega^{\star}$ symmetrization, the proofs can be done is the same spirit as their Euclidean counterparts, and no change of variables is required.
2.3. An iterated Talenti’s principle
The following result is a Talenti’s principle on Riemannian manifolds supporting the isoperimetric inequality (1). The statement is formulated using the notations of the previous section. The proof is analogous to the main result of Chen and Li [6, Theorem 1.1], which expoits the isoperimetric inequality (4).
Theorem 2.2.
Let $(M,g)$ is complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a bounded domain with smooth boundary and let $\Omega^{\star}$ denote its symmetrization. Suppose that $u$ and $v$ are weak solutions of the problems
$\begin{cases}-\Delta_{g}u=f,&\text{in }\Omega,\\ u=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}-\Delta_{g_{0}}v=f^{\star},&\text{in }\Omega^{\star},\\ v=0,&\text{on }\partial\Omega^{\star},\end{cases}$ (12)
respectively. Then for almost every $x\in\Omega^{\star}$ one has
$(c_{g})^{2}\cdot u^{\star}(x)\leq v(x).$ (13)
Next, in the spirit of Gazzola, Grunau, and Sweers [11], we establish an iterated version of Theorem 2.2, which enables us to prove higher-order Sobolev and Rellich inequalities. Our result reads as follows.
Theorem 2.3.
Let $(M,g)$ be a complete, non-compact Riemannian manifold supporting inequality (1). Assume that $\Omega\subset M$ is a bounded domain with smooth boundary and let $\Omega^{\star}$ denote its symmetrization. Let $k\geq 1$ be an integer, and suppose that $u$ and $v$ are weak solutions of the problems
$\begin{cases}(-\Delta_{g})^{k}u=f,&\text{in }\Omega,\\ u=\Delta_{g}u=\dots=\Delta_{g}^{k-1}u=0,&\text{on }\partial\Omega,\end{cases}$ (14)
and
$\begin{cases}(-\Delta_{g_{0}})^{k}v=f^{\star},&\text{in }\Omega^{\star},\\ v=\Delta_{g_{0}}v=\dots=\Delta_{g_{0}}^{k-1}v=0,&\text{on }\partial\Omega^{\star}.\end{cases}$ (15)
Then for almost every $x\in\Omega^{\star}$ one has
$(c_{g})^{2k}\cdot u^{\star}(x)\leq v(x).$ (16)
Proof.
We proceed by finite induction. When $k=1$, Theorem 2.3 follows from Theorem 2.2. In the sequel, we prove Theorem 2.3 for an arbitrary $k\geq 2$, assuming that it is true for k1k-1. Let
$\widetilde{u}\stackrel{{\scriptstyle\textup{def}}}{{=}}-\Delta_{g}u\quad\text{and}\quad\widetilde{v}\stackrel{{\scriptstyle\textup{def}}}{{=}}-\Delta_{g_{0}}v.$
On the one hand, according to problems (14) & (15), one has
$\begin{cases}(-\Delta_{g})^{k-1}\widetilde{u}=f,&\text{in }\Omega,\\ \widetilde{u}=\Delta_{g}\widetilde{u}=\dots=\Delta_{g}^{k-2}\widetilde{u}=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}(-\Delta_{g_{0}}\widetilde{v})^{k-1}=f^{\star},&\text{in }\Omega^{\star},\\ \widetilde{v}=\Delta_{g_{0}}\widetilde{v}=\dots=\Delta_{g_{0}}^{k-1}\widetilde{v}=0,&\text{on }\partial\Omega^{\star}.\end{cases}$
Hence, our assumption (Theorem 2.3 is true for k1k-1) implies that for a.e. $x\in\Omega$ one has
$(c_{g})^{2(k-1)}\cdot\widetilde{u}^{\star}(x)\leq\widetilde{v}(x).$
On the other hand, we also have $\widetilde{u}=0$ and $\widetilde{v}=0$ on $\partial\Omega$ and $\partial\Omega^{\star}$, respectively, thus we obtain
$\begin{cases}-\Delta_{g}u=\widetilde{u}&\text{in }\Omega,\\ u=0,&\text{on }\partial\Omega,\end{cases}\quad\text{and}\quad\begin{cases}-\Delta_{g_{0}}v=\widetilde{v}&\text{in }\Omega^{\star},\\ v=0,&\text{on }\partial\Omega^{\star}.\end{cases}$
Thus, combining the Euclidean maximum principle for $\Delta$ with another application of Theorem 2.2, implies the inequality (16).
Remark 2.4.
We note that the non-iterated version of Talenti’s principle is available on Finsler manifolds with non-negative weighted Ricci curvature (see Mester [18]) and on the so-called $\mathsf{RCD}(0,N)$ spaces (see Wu [21]). Depending on the way of symmetrization (cf. Remark 2.1), inequality (13) may also be written in the form $u^{\sharp}\leq v$, which is the case in [6] & [18]. For simplicity, our results are presented on Riemannian manifolds; however, by following their line of thought, one may extend them to the aforementioned settings, taking into account the technicalities of corresponding spaces.
3. Sobolev inequalities
The goal of this section is twofold. First, we prove Theorem 1.1, that is, we establish higher-order Sobolev inequalities on Riemannian manifolds supporting inequality (1). Next, we present a sharpness result concerning a particular Sobolev inequality, see Proposition 3.1.
Proof of Theorem 1.1.
Since $u\in C_{0}^{\infty}(M)$, there exists a bounded domain $\Omega\in M$, such that $u$ (and its derivatives) vanish outside of $\Omega$. Thus, it is enough to prove the Sobolev inequality (2) on an arbitrary bounded domain $\Omega\subset M$. Without loss of generality, one may assume that $u>0$. We distinguish two cases.
Case 1: $m=2k$. Let $f\colon\Omega\to\mathbb{R}$ defined by
$f=(-\Delta_{g}^{k})u.$ (17)
If follows that, $u$ is a solution of problem (14), and let $v$ be the solution of the symmetrized problem (15). By the iterated Talenti’s principle (see Theorem 2.3) one has for almost every $x\in\Omega$ that
$(c_{g})^{2k}\cdot u^{\star}(x)\leq v(x),$ (18)
First, relations (14) & (15) and the Cavalieri principle (see relation (9)) implies
$\int_{\Omega}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}=\int_{\Omega}|\Delta_{g}^{k}u|^{p}\,\mathrm{d}v_{g}=\int_{\Omega}|f|^{p}\,\mathrm{d}v_{g}=\int_{\Omega^{\star}}(f^{\star})^{p}\,\mathrm{d}v_{g_{0}}=\int_{\Omega^{\star}}|\Delta_{g_{0}}^{k}v|^{p}\,\mathrm{d}v_{g_{0}}.$
Next, the Euclidean Sobolev inequality and relation (18) yields
$\int_{\Omega^{\star}}|\Delta_{g_{0}}^{k}v|^{p}\,\mathrm{d}v_{g_{0}}\geq S_{m,p}\cdot\left(\int_{\Omega^{\star}}|v|^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}\geq(c_{g})^{2kp}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}(u^{\star})^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}.$
Finally, using the relation $2k=m$ and the Cavalieri principle leads us to
$(c_{g})^{2kp}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}(u^{\star})^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}u^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
which is precisely inequality (2).
Case 2: $m=2k+1$. Define $f$ as in relation (17), then $u$ is a solution of the problem (14) and let $v$ be a solution of the symmetrized problem (15). Note that, here we use the same problem as in the previous case (characterized by the parameter $k$).
On the one hand, using relations (14) & (15) and the Pólya–Szegő inequality (see relation (10)) we obtain
$\int_{\Omega}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}$ $=\int_{\Omega}|\nabla_{g}(\Delta_{g}^{k}u)|^{p}\,\mathrm{d}v_{g}=\int_{\Omega}|\nabla_{g}f|^{p}\,\mathrm{d}v_{g}$
$\geq(c_{g})^{p}\cdot\int_{\Omega^{\star}}(\nabla_{g_{0}}f^{\star})^{p}\,\mathrm{d}v_{g_{0}}=(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}(\Delta_{g_{0}}^{k}v)|^{p}\,\mathrm{d}v_{g_{0}}.$
On the other hand, the Euclidean Sobolev inequality, relation (18), and the Cavalieri principle yields
$(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}(\Delta_{g_{0}}^{k}v)|^{p}\,\mathrm{d}v_{g_{0}}$ $\geq(c_{g})^{p}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}|v|^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}$
$\geq(c_{g})^{(2k+1)p}\cdot S_{m,p}\cdot\left(\int_{\Omega^{\star}}(u^{\star})^{p^{*}}\,\mathrm{d}v_{g_{0}}\right)^{\frac{p}{p^{*}}}$
$=(c_{g})^{mp}\cdot S_{m,p}\cdot\left(\int_{\Omega}u^{p^{*}}\,\mathrm{d}v_{g}\right)^{\frac{p}{p^{*}}},$
which ends the proof. ∎
We conclude this section with a sharpness result concerning the Sobolev inequality (2) on manifolds with $\mathbf{Ric}_{g}\geq 0$ and Euclidean volume growth, in the particular case when $m=2$ and $p=2$. This finding extends a result of Barbosa and Kristály [3]: If a second-order Sobolev inequality with the optimal Euclidean constant holds on a Riemannian manifold $(M,g)$ with non-negative Ricci curvature, then $(M,g)$ is isometric to the Euclidean space $(\mathbb{R}^{n},g_{0})$.
Proposition 3.1.
Let $(M,g)$ be a complete, non-compact $n$-dimensional Riemannian manifold, with $n>4$, $\mathbf{Ric}_{g}\geq 0$, and Euclidean volume growth. Let $\rho(\cdot)=d_{g}(x_{0},\cdot)$ for some $x_{0}\in M$ and suppose that $\rho\Delta_{g}\rho\geq n-5$. Then the inequality
$\int_{M}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq\mathsf{AVR}_{g}^{\frac{4}{n}}\cdot S_{2,2}\cdot\left(\int_{M}|u|^{\frac{2n}{n-4}}\,\mathrm{d}v_{g}\right)^{\frac{n-4}{n}},\quad\forall u\in C_{0}^{\infty}(M),$ (19)
is sharp in the sense that the constant $\mathsf{AVR}_{g}^{\frac{4}{n}}\cdot S_{2,2}$ cannot be improved.
Proof.
The proof can be done in the spirit of Kristály [16, Proposition 2.3 & Theorem 1.1] as follows.
Suppose that there exists a constant $C>\mathsf{AVR}_{g}^{\frac{4}{n}}\cdot S_{2,2}$ such that for every $u\in C_{0}^{\infty}(M)$ one has
$\int_{M}(\Delta_{g}u)^{2}\,\mathrm{d}v_{g}\geq C\cdot\left(\int_{M}|u|^{\frac{2n}{n-4}}\,\mathrm{d}v_{g}\right)^{\frac{n-4}{n}}.$ (20)
To derive a contradiction, we use the Talentian bubbles defined by
$u_{\lambda}\stackrel{{\scriptstyle\textup{def}}}{{=}}t_{\lambda}(\rho),\quad\text{where}\quad t_{\lambda}(r)\stackrel{{\scriptstyle\textup{def}}}{{=}}(\lambda+r^{2})^{\frac{4-n}{2}},\quad\lambda,r>0,$
as test functions in inequality (20); this is allowed, since they can be be approximated by $C_{0}^{\infty}(M)$ functions.
On the one hand, using the chain rule and the eikonal equation ($|\nabla_{g}\rho|=1$ a.e. in $M$), for almost every $x\in M$ we obtain
$(\Delta_{g}u_{\lambda})^{2}=(n-4)^{2}(\lambda+\rho^{2})^{-n}\left[\lambda+(3-n)\rho^{2}+(\lambda+\rho^{2})\rho\Delta_{g}\rho\right]^{2}.$ (21)
On the other hand, by assumption and the Laplacian comparison theorem applied for the case $\mathbf{Ric}_{g}\geq 0$ (see e.g. Gallot, Hulin, and Lafontaine [9, Theorem 3.101]) we also have
$n-1\geq\rho\Delta_{g}\rho\geq n-5,$
which yields
$\left[\lambda+(3-n)\rho^{2}+(\lambda+\rho^{2})\rho\Delta_{g}\rho\right]^{2}\leq(n\lambda+2\rho^{2})^{2}.$ (22)
According to the relations (20), (21), and (22), for every $\lambda>0$ we obtain
$(n-4)^{2}\int_{M}(\lambda+\rho^{2})^{-n}(n\lambda+2\rho^{2})^{2}\,\mathrm{d}v_{g}\geq C\cdot\left(\int_{M}(\lambda+\rho^{2})^{-n}\,\mathrm{d}v_{g}\right)^{\frac{n-4}{n}}.$
Multiplying by $\lambda^{\frac{n-4}{2}}$ and using the notation
$\mathcal{H}(\lambda,s)\stackrel{{\scriptstyle\textup{def}}}{{=}}\int_{M}(\lambda+d_{g}^{2})^{-n}\,\mathrm{d}v_{g},\quad\lambda,s>0,$
leads us to
$(n-4)^{2}\left(4\lambda^{\frac{n}{2}-2}\mathcal{H}(\lambda,n-2)+4(n-2)\lambda^{\frac{n}{2}-1}\mathcal{H}(\lambda,n-1)+(n-2)^{2}\lambda^{\frac{n}{2}}\mathcal{H}(\lambda,n)\right)\geq C\cdot\left(\lambda^{\frac{n}{2}}\mathcal{H}(\lambda,n)\right)^{\frac{n-4}{n}}.$ (23)
As a final step we take the limit as $\lambda\to\infty$ in inequality (23). To do this, we use the following asymptotic property of $\mathcal{H}$ for $s>n/2$:
$\lim_{\lambda\to\infty}\lambda^{s-\frac{n}{2}}\mathcal{H}(\lambda,s)=\omega_{n}\cdot\mathsf{AVR}_{g}\cdot\frac{\Gamma\left(\frac{n}{2}+1\right)\Gamma\left(s-\frac{n}{2}\right)}{\Gamma(s)},$
(see e.g. Kristály [16, Proposition 2.3/(ii)]) and the following property of the gamma function:
$\Gamma(z+1)=z\Gamma(z),\quad\forall z>0.$
After taking the limit and simplifying the left hand side, we obtain
$(n-4)(n-2)n(n+2)\frac{\Gamma\!\left(\frac{n}{2}\right)}{\Gamma(n)}$ $\geq\frac{C}{\pi^{2}}\mathsf{AVR}_{g}^{-\frac{4}{n}}\left(\frac{\Gamma\left(\frac{n}{2}\right)}{\Gamma(n)}\right)^{\frac{n-4}{n}}.$
Observe that the letter inequality is equivalent to
$S_{2,2}\cdot\mathsf{AVR}_{g}^{\frac{4}{n}}\geq C,$
which is a contradiction. This completes the proof. ∎
Remark 3.2.
One may wonder whether similar (possibly higher order) assumptions as $\rho\Delta_{g}\rho\geq n-5$ yield the sharpness of higher-order Sobolev inequalities. An important ingredient of the previous proof is to guarantee inequality (22), which can be rewritten as
$(\Delta_{g}u_{\lambda})^{2}\leq L(\rho)^{2},\quad\text{where}\quad L(|x|)=\Delta_{g_{0}}t_{\lambda}(|x|).$
It turns out that the higher-order version of the above pointwise inequality implies rigidity (the manifold needs to be Euclidean). Thus, a possible proof of the sharpness (if exists) may rely on non-pointwise estimates.
4. Rellich inequalities and improvements
In this section we prove Theorem 1.2, that is we establish higher-order Rellich inequalities on Riemannian manifolds supporting inequality (1). After that, we present various improved Rellich inequalities, which can be proved in a similar manner.
Proof of Theorem 1.2.
The proof follows a similar line of thoughts as the proof of Theorem 1.1: We observe that it is enough to prove it for smooth positive functions $u$ on bounded domains $\Omega$. In addition, we define $f=(-\Delta_{g})^{k}u$ and consider the problem (14), and its symmetrization (15), whose solution is denoted by $v$.
On the one hand, if $m=2k$, by applying the Cavalieri principle (see (9)), the iterated Talenti’s principle (see Theorem 2.3), the Euclidean Rellich inequality, and the Hardy–Littlewood inequality (see (11)) we get
$\int_{\Omega}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}$ $=\int_{\Omega}|\Delta_{g}^{k}u|^{p}\,\mathrm{d}v_{g}=\int_{\Omega}|f|^{p}\,\mathrm{d}v_{g}=\int_{\Omega^{\star}}(f^{\star})^{p}\,\mathrm{d}v_{g_{0}}=\int_{\Omega^{\star}}|\Delta_{g_{0}}^{k}v|^{p}\,\mathrm{d}v_{g_{0}}$
$\geq R_{m,p}\cdot\int_{\Omega^{\star}}\frac{|v|^{p}}{|x|^{mp}}\,\mathrm{d}v_{g_{0}}\geq(c_{g})^{2kp}\cdot R_{m,p}\cdot\int_{\Omega^{\star}}\frac{(u^{\star})^{p}}{|x|^{mp}}\,\mathrm{d}v_{g_{0}}$
$\geq(c_{g})^{mp}\cdot R_{m,p}\cdot\int_{\Omega}\frac{|u|^{p}}{\rho^{mp}}\,\mathrm{d}v_{g}.$
On the other hand, if $m=2k+1$, we also use the Pólya–Szegő inequality (see (10)) to obtain
$\int_{\Omega}|D_{g}^{m}u|^{p}\,\mathrm{d}v_{g}$ $=\int_{\Omega}|\nabla_{g}\Delta_{g}^{k}u|^{p}\,\mathrm{d}v_{g}=\int_{\Omega}|\nabla_{g}f|^{p}\,\mathrm{d}v_{g}\geq(c_{g})^{p}\cdot\int_{\Omega^{\star}}(\nabla_{g_{0}}f^{\star})^{p}\,\mathrm{d}v_{g_{0}}$
$=(c_{g})^{p}\cdot\int_{\Omega^{\star}}|\nabla_{g_{0}}\Delta_{g_{0}}^{k}v|^{p}\,\mathrm{d}v_{g_{0}}\geq(c_{g})^{p}\cdot R_{m,p}\cdot\int_{\Omega^{\star}}\frac{|v|^{p}}{|x|^{mp}}\,\mathrm{d}v_{g_{0}}$
$\geq(c_{g})^{(2k+1)p}\cdot R_{m,p}\cdot\int_{\Omega^{\star}}\frac{(u^{\star})^{p}}{|x|^{mp}}\,\mathrm{d}v_{g_{0}}\geq(c_{g})^{mp}\cdot R_{m,p}\cdot\int_{\Omega}\frac{|u|^{p}}{\rho^{mp}}\,\mathrm{d}v_{g},$
which concludes the proof. ∎
Proof of Theorem 1.3.
Observe that the proof of Theorem 1.2 provides a ‘recipe’ for threating improved Rellich inequalities of the form
$\int_{\Omega^{\star}}|D_{g_{0}}^{m}u|^{p}\,\mathrm{d}v_{g_{0}}\geq C_{m,p}\cdot\int_{\Omega^{\star}}|u|^{p}h(|x|)\,\mathrm{d}v_{g_{0}},\quad u\in C_{0}^{\infty}(\Omega^{\star}).$
The only difference is that we need to use the Hardy–Littlewood inequality (see (11)) for a generic $h$ instead of $t\mapsto\frac{1}{t^{mp}}$. ∎
In the rest of the paper, we list various applications of Theorem 1.3. For brevity, in each case, first we indicate a reference for the Euclidean version, next we state the Riemannian extension of the corresponding inequality.
  • Adimurthi, Grossi, and Santrac [1, Theorem 2.1/(a)]: For any function $h$ let $h_{[0]}(t)=t$, $h_{[1]}(t)=h(t)$ and $h_{[s]}(t)=h(h_{[s-1]}(t))$, for all $s\geq 2$. Define $\ell(t)=\frac{1}{1-\log(t)}$, and let $T\geq 1$ be an integer and $R>0$. Then for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{T}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\prod_{s=1}^{j}\ell_{[s]}^{2}\left(\frac{|x|}{R}\right)\,\mathrm{d}v_{g}\right).$
  • Ghoussoub and Moradifam [12, Corollary 3.8/1]: Let $n\geq 5$, $\Omega\subseteq M$ be bounded, $k\geq 1$ be an integer, $R$ denote the radius of $\Omega^{\star}$ and define $r=R\cdot\exp_{[k-1]}(e)$. Then for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{k}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\left(\prod_{i=1}^{j}\log_{[i]}\left(\frac{r}{|x|}\right)\right)^{-2}\,\mathrm{d}v_{g}\right).$
  • Gazzola, Grunau, and Mitidieri [10, Theorem 2]: Let $p\geq 2$, $n\geq 2p$, and $\Omega\subset M$ be bounded, then for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|\Delta_{g}u|^{p}\,\mathrm{d}v_{g}$ $\geq(c_{g})^{2p}\cdot\left(\frac{(n-2p)(p-1)n}{p^{2}}\right)^{p}\cdot\int_{\Omega}\frac{|u|^{p}}{\rho^{2p}}\,\mathrm{d}v_{g}$
    $\qquad+(c_{g})^{2p}\cdot\frac{4(p-1)^{p}(n-2p)^{p-1}n^{p-1}\gamma}{p^{2p-1}}\cdot\left(\frac{\omega_{n}}{\operatorname{vol}_{g}(\Omega)}\right)^{\frac{2}{n}}\cdot\int_{\Omega}\frac{|u|^{p}}{\rho^{2p-2}}\,\mathrm{d}v_{g}$
    $\qquad+(c_{g})^{2p}\cdot\Gamma\cdot\left(\frac{\omega_{n}}{\operatorname{vol}_{g}(\Omega)}\right)^{\frac{2p}{n}}\cdot\int_{\Omega}|u|^{p}\,\mathrm{d}v_{g}.$
    The constants appearing in the above inequality are defined as follows. Let $\vartheta=4+\frac{n(p-2)}{p}$, introduce the function space $X=\{v\in C^{2}([0,1]):v^{\prime}(0)=v(1)=0,v\not\equiv 0\}$, and denote $\Delta_{\vartheta}v(r)=v^{\prime\prime}(r)+\frac{\vartheta-1}{r}v^{\prime}(r)$. Define
    $\gamma\stackrel{{\scriptstyle\textup{def}}}{{=}}\displaystyle\inf_{X}\frac{\int_{0}^{1}r|v^{\prime}(r)|^{2}\,\mathrm{d}r}{\int_{0}^{1}rv(r)^{2}\,\mathrm{d}r}\quad\text{and}\quad\Gamma\stackrel{{\scriptstyle\textup{def}}}{{=}}\displaystyle\max\left\{\frac{(p-1)^{p-1}(n-2p)^{p-2}n^{p-2}}{p^{2(p-2)}}\Lambda_{\vartheta},\lambda_{\vartheta}\right\},$
    where $\lambda_{\vartheta}\stackrel{{\scriptstyle\textup{def}}}{{=}}\displaystyle\inf_{X}\frac{\int_{0}^{1}r^{2p-1}|\Delta_{\vartheta}v(r)|^{p}\,\mathrm{d}r}{\int_{0}^{1}r^{2p-1}|v(r)|^{p}\,\mathrm{d}r}$ and $\Lambda_{\vartheta}\stackrel{{\scriptstyle\textup{def}}}{{=}}\displaystyle\inf_{X}\frac{\int_{0}^{1}r^{3}|v(r)|^{p-2}(\Delta_{\vartheta}v(r))^{2}\,\mathrm{d}r}{\int_{0}^{1}r^{2p-1}|v(r)|^{p}\,\mathrm{d}r}$.
    In particular, when $p=2$, we obtain the following second order Brezis–Vazquez-type inequality:
    $\int_{\Omega}(\Delta_{g}u)^{2}\,\mathrm{d}v_{g}$ $\geq(c_{g})^{4}\cdot\frac{n^{2}(n-4)^{2}}{16}\cdot\int_{\Omega}\frac{u^{2}}{\rho^{4}}\,\mathrm{d}v_{g}+(c_{g})^{4}\cdot\frac{n(n-4)}{2}\cdot\Lambda(2)\cdot\left(\frac{\omega_{n}}{\operatorname{vol}_{g}(\Omega)}\right)^{\frac{2}{n}}\cdot\int_{B}\frac{u^{2}}{\rho^{2}}\,\mathrm{d}v_{g}$
    $\qquad+(c_{g})^{4}\cdot{\Lambda(4)^{2}}\cdot\left(\frac{\omega_{n}}{\operatorname{vol}_{g}(\Omega)}\right)^{\frac{4}{n}}\cdot\int_{\Omega}u^{2}\,\mathrm{d}v_{g},\quad\forall u\in C_{0}^{\infty}(\Omega),$
    where $\Lambda(n)=\displaystyle\inf_{X}\frac{\int_{0}^{1}r^{n-1}|v^{\prime}(r)|^{2}\,\mathrm{d}r}{\int_{0}^{1}r^{n-1}v(r)^{2}\,\mathrm{d}r}$.
  • Gazzola, Grunau, and Mitidieri [10, Corollary 2]: Let $\Omega\subset M$ be bounded and $m\geq 1$ be an integer with $n>2m$. Then there exist some constants $c_{1},c_{2},\dots,c_{m}$ depending only on $m$, $n$, and $\operatorname{vol}_{g}(\Omega)$ such that for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|D^{m}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{2m}\cdot\left(L_{m}\cdot\int_{\Omega}\frac{u^{2}}{\rho^{2m}}\,\mathrm{d}v_{g}+\sum_{s=1}^{m}c_{s}\int_{\Omega}\frac{u^{2}}{\rho^{2(m-s)}}\,\mathrm{d}v_{g}\right),$
    where $L_{m}=\displaystyle\frac{1}{4^{m}}\prod_{s=1}^{m}(n+2m-4j)^{2}$.